her three test scores were three consecutive integers. if the sum of her scores was 285, determined her three scores
Answer by TimothyLamb(4379)

k = j + 1
j = i + 1
i + j + k = 285
---
i + j + j + 1 = 285
i + 2j + 1 = 285
i + 2(i + 1) + 1 = 285
i + 2i + 2 + 1 = 285
3i + 3 = 285
3i = 282
---
i = 94
j = 95
k = 96
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php